OpenStudy (anonymous):

A charity fair raised $6,000 by selling 500 lottery tickets. There were two types of lottery tickets: A tickets cost $10 each, and B tickets cost $60 each. How many tickets of each type were sold? A)250, 250 B)315, 185 C)370, 230 D)410, 90 E)480, 20

OpenStudy (anonymous):

the answer is C

OpenStudy (anonymous):

So the equations are 10a + 60b = 6000 and a + b = 500 You could just solve those equations, but I would guess and check on a multiple choice test. 6000/500 = 12, which is so close to 10 that a large majority of the tickets sold were 'a' tickets. 10(480) + 60(20) = 4800 + 1200 = 6000 bingo. The answer is c. 480, 20

OpenStudy (anonymous):

@Kemist thats what i got too thanxx

OpenStudy (anonymous):

You're whalecummm :DD

OpenStudy (anonymous):

@Kemist its really E ... if its 480,20 ^^ )

OpenStudy (anonymous):

Just copied it from a previous question. But atleast you got the aswer.

OpenStudy (anonymous):

lol yup thanx bunches :3